# Sequential Randomization Tests Using e-values: Applications for trial monitoring

### v4

@@ 1 Introduction
Monitoring clinical trials, specially in the context of acutely ill patients, is of paramount importance. Traditional methods include $\alpha-spending$ functions. When interim analyses are spaced, evidence grows unnoticed between the interim analyses. This may both delay the implementation of potentially useful strategies or prolong trials when safety signals arise early.
@@ 4.2 Betting strategy for continuous outcomes
    A center and scale for the outcomes observed so far. Outcomes con be all all over the place. We need to standardize.
@@ 4.6 Simulation overview
A visual representation can be found in figure below.
@@ 6 Discussion
The e-Rt is an nonparametric sequential test for randomized trials based on the betting framework for e-values (i-bet pmlr-v177-duan22a). The method requires only that treatment assignment is randomized — no distributional assumptions about outcomes are needed. This makes it a robust complement to model-based analyses. It can be extended to continuous endpoints and time-to-event analyses.
@@ 7.4 Version Control
    Forth Version (This version): Correction on signal direction on -RTC. Text adjustments
@@ Appendix A R Code
# e-RT Simulations: Type I Error and Power (Unified)
@@
if (outcome[i] == 1) {
lambda <- 0.5 + 0.5 * c_i * delta_hat
@@
if (treatment[i] == 1) {
multiplier <- lambda / p
} else {
multiplier <- (1 - lambda) / (1 - p)
@@
return(data.frame(treatment = treatment, outcome = outcome))
# Unified simulation function
simulate_eRT <- function(n_sims = 5000,
p_ctrl,
p_trt = NULL,
hypothesized_ARR = NULL,
target_power = 0.80,
alpha = 0.05,
burn_in = 50,
ramp = 100) {
# Determine hypothesized ARR for sample size calculation
@@
if (is.null(hypothesized_ARR)) {
hypothesized_ARR <- p_ctrl - p_trt
# Calculate sample size based on hypothesized effect
@@
sim_type <- "Power"
[…]
p_ctrl, hypothesized_ARR, target_power * 100))
[…]
p_ctrl, p_trt, true_ARR))
[…]
@@
wealth <- compute_eRT(trial\$treatment, trial\$outcome, burn_in = burn_in, ramp = ramp)
@@
median_stop <- median(first_crossing, na.rm = TRUE)
[…]
if (sim_type == "Power") {
[…]
} else {
[…]
return(list(
sim_type = sim_type,
@@
target_power = target_power,
alpha = alpha,
median_crossing = median_stop,
@@
# Plot trajectories
plot_trajectories <- function(n_trials = 30,
p_ctrl,
p_trt = NULL,
hypothesized_ARR = NULL,
target_power = 0.80,
alpha = 0.05,
burn_in = 50,
ramp = 100,
@@
if (is.null(hypothesized_ARR)) {
hypothesized_ARR <- p_ctrl - p_trt
if (is.null(p_trt)) {
p_trt <- p_ctrl # Null
# Calculate sample size
@@
if (true_ARR == 0) {
title <- sprintf("e-RTUnderNull(n=%d,designedfor%.0f%%ARR)",
@@
title <- sprintf("e-RTUnderAlternative(n=%d,%.0f%%power)",
@@
trajectories <- list()
for (i in 1:n_trials) {
@@
wealth <- compute_eRT(trial\$treatment, trial\$outcome, burn_in = burn_in, ramp = ramp)
trajectories[[i]] <- data.frame(
patient = 1:n_patients,
wealth = wealth,
trial = i
@@
p <- ggplot(df, aes(x = patient, y = wealth, group = trial)) +
@@
labs(
title = title,
@@
x = "Patient",
y = "e-value(logscale)"
@@
theme(
plot.title = element_text(face = "bold"),
panel.grid.minor = element_blank()
return(p)
# Run all simulations
run_all <- function(n_sims = 5000,
p_ctrl = 0.40,
ARRs = c(0.05, 0.10),
target_powers = c(0.80, 0.90),
alpha = 0.05,
seed = 42) {
set.seed(seed)
results <- list()
# Create grid of scenarios
scenarios <- expand.grid(
hypothesized_ARR = ARRs,
target_power = target_powers,
stringsAsFactors = FALSE
# For each scenario, run Type I error and Power
all_results <- data.frame()
for (i in 1:nrow(scenarios)) {
hyp_ARR <- scenarios\$hypothesized_ARR[i]
tgt_pow <- scenarios\$target_power[i]
[…]
hyp_ARR * 100, tgt_pow * 100))
[…]
# Type I error (null true, same sample size)
[…]
t1 <- simulate_eRT(
n_sims = n_sims,
p_ctrl = p_ctrl,
p_trt = NULL, # Null is true
hypothesized_ARR = hyp_ARR,
target_power = tgt_pow,
alpha = alpha
# Power (alternative true)
[…]
pow <- simulate_eRT(
n_sims = n_sims,
p_ctrl = p_ctrl,
p_trt = p_ctrl - hyp_ARR, # True effect = hypothesized
target_power = tgt_pow,
alpha = alpha
all_results <- rbind(all_results, data.frame(
hypothesized_ARR = hyp_ARR,
target_power = tgt_pow,
n_patients = t1\$n_patients,
type1_error = t1\$rejection_rate,
type1_se = t1\$se,
eRT_power = pow\$rejection_rate,
power_se = pow\$se,
median_crossing = pow\$median_crossing
results\$summary <- all_results
# Print summary
[…]
print(all_results %>%
mutate(
hypothesized_ARR = sprintf("%.0f%%", hypothesized_ARR * 100),
target_power = sprintf("%.0f%%", target_power * 100),
type1_error = sprintf("%.3f", type1_error),
eRT_power = sprintf("%.1f%%", eRT_power * 100)
select(hypothesized_ARR, target_power, n_patients,
type1_error, eRT_power, median_crossing))
return(results)
# Run
if (interactive()) {
results <- run_all(
n_sims = 5000,
p_ctrl = 0.40,
ARRs = c(0.05, 0.10),
target_powers = c(0.80, 0.90),
alpha = 0.05
# Trajectory plots - Alternative
p1 <- plot_trajectories(p_ctrl = 0.40, p_trt = 0.30, target_power = 0.80)
ggsave("traj_alt_10pct_80pow.pdf", p1, width = 8, height = 5)
p2 <- plot_trajectories(p_ctrl = 0.40, p_trt = 0.30, target_power = 0.90)
ggsave("traj_alt_10pct_90pow.pdf", p2, width = 8, height = 5)
# Trajectory plots - Null
p3 <- plot_trajectories(p_ctrl = 0.40, p_trt = NULL, hypothesized_ARR = 0.10, target_power = 0.80)
ggsave("traj_null_10pct_80pow.pdf", p3, width = 8, height = 5)
p4 <- plot_trajectories(p_ctrl = 0.40, p_trt = NULL, hypothesized_ARR = 0.10, target_power = 0.90)
ggsave("traj_null_10pct_90pow.pdf", p4, width = 8, height = 5)
# e-RTC (Randomization E-Process for Continuous Outcomes)
# Type I Error and Power Simulations (Unified)
setwd("")
suppressPackageStartupMessages({
library(tidyverse)
# Core e-RTC function
# treatment: 0/1 vector (0 = control, 1 = intervention)
# outcome : continuous outcome (numeric)
# p : allocation prob for treatment (default 0.5)
# burn_in : number of patients before bets start ramping up
# ramp : additional patients over which betting strength ramps to max
# c_max : maximum betting strength (0 < c_max <= 0.5 recommended)
compute_eRTC <- function(treatment, outcome,
p = 0.5,
burn_in = 20,
ramp = 50,
c_max = 0.6) {
@@
# Require enough past data before betting
@@
g_i <- s_i / (1 + abs(s_i)) # in (-1,1), robust to outliers
# Learn direction from past data (NEW)
@@
# Handle edge cases early in trial
if (is.nan(mean_trt) $|$ is.nan(mean_ctrl)) {
direction <- 0
} else {
direction <- sign(mean_trt - mean_ctrl)
# Ramping betting strength
@@
# Wealth update based on actual assignment
if (treatment[i] == 1) {
multiplier <- lambda / p
} else {
multiplier <- (1 - lambda) / (1 - p)
@@
# Simulate a single trial with continuous outcomes
# n : total number of patients
# p_trt : probability of being assigned to treatment (default 0.5)
# mu_ctrl : mean in control arm
# mu_trt : mean in treatment arm
# sd : common standard deviation (default 1)
simulate_trial_continuous <- function(n,
p_trt = 0.5,
mu_ctrl = 0,
mu_trt = 0,
sd = 1) {
@@
return(data.frame(treatment = treatment, outcome = outcome))
# Unified simulation function for e-RTC
# n_sims : number of simulated trials
# mu_ctrl : control mean (default 0)
# mu_trt : treatment mean (NULL => null, mu_trt = mu_ctrl)
# hypothesized_d: Cohen’s d used for sample size calc (delta/sd)
# target_power : design power (0.80, 0.90)
# alpha : significance level
# burn_in, ramp, c_max: e-RTC tuning parameters
simulate_eRTC <- function(n_sims = 5000,
mu_ctrl = 0,
mu_trt = NULL,
hypothesized_d,
target_power = 0.80,
alpha = 0.05,
burn_in = 20,
ramp = 50,
c_max = 0.6) {
if (missing(hypothesized_d)) {
stop("Mustspecifyhypothesized_d(Cohen’sdfordesign).")
# Treat SD = 1 for design (Cohen’s d = delta / sd)
@@
# Sample size via t-test (two-sample, equal sizes)
$s$ <- power.t.test(delta = delta_design,
sd = sd_true,
power = target_power,
sig.level = alpha,
type = "two.sample",
alternative = "two.sided")
@@
sim_type <- "Power"
[…]
mu_ctrl, hypothesized_d, target_power * 100))
[…]
mu_ctrl, mu_trt, true_d))
[…]
@@
trial <- simulate_trial_continuous(n_patients,
mu_ctrl = mu_ctrl,
mu_trt = mu_trt,
sd = sd_true)
wealth <- compute_eRTC(trial\$treatment,
trial\$outcome,
p = 0.5,
burn_in = burn_in,
ramp = ramp,
c_max = c_max)
@@
median_stop <- median(first_crossing, na.rm = TRUE)
[…]
rejection_rate - 1.96 * se,
rejection_rate + 1.96 * se))
if (sim_type == "Power") {
[…]
} else {
[…]
return(list(
sim_type = sim_type,
@@
target_power = target_power,
alpha = alpha,
median_crossing = median_stop,
@@
# Plot trajectories under null / alternative
# n_trials : number of trajectories to overlay
# mu_ctrl, mu_trt, hypothesized_d, target_power, alpha, burn_in, ramp, c_max
plot_trajectories_eRTC <- function(n_trials = 30,
mu_ctrl = 0,
mu_trt = NULL,
hypothesized_d,
target_power = 0.80,
alpha = 0.05,
burn_in = 20,
ramp = 50,
c_max = 0.6,
title = NULL) {
if (missing(hypothesized_d)) {
stop("Mustspecifyhypothesized_d(Cohen’sd).")
@@
delta_design <- hypothesized_d * sd_true
$s$ <- power.t.test(delta = delta_design,
sd = sd_true,
power = target_power,
sig.level = alpha,
type = "two.sample",
alternative = "two.sided")
@@
if (true_d == 0) {
title <- sprintf(
"e-RTCUnderNull(n=%d,designedford=%.2f)",
n_patients, hypothesized_d
@@
title <- sprintf(
"e-RTCUnderAlternative(n=%d,trued=%.2f,%.0f%%powerdesign)",
n_patients, true_d, target_power * 100
trajectories <- vector("list", n_trials)
for (i in 1:n_trials) {
trial <- simulate_trial_continuous(n_patients,
mu_ctrl = mu_ctrl,
mu_trt = mu_trt,
sd = sd_true)
wealth <- compute_eRTC(trial\$treatment,
trial\$outcome,
p = 0.5,
burn_in = burn_in,
ramp = ramp,
c_max = c_max)
trajectories[[i]] <- data.frame(
patient = 1:n_patients,
wealth = wealth,
trial = i
@@
p <- ggplot(df, aes(x = patient, y = wealth, group = trial)) +
@@
labs(
title = title,
@@
x = "Patient",
y = "e-value(logscale)"
annotate("text",
x = n_patients * 0.95,
y = (1 / alpha) * 1.5,
label = sprintf("1/alpha=%.0f", 1 / alpha),
color = "red", hjust = 1) +
@@
theme(
plot.title = element_text(face = "bold"),
panel.grid.minor = element_blank()
return(p)
# Run all scenarios (grid over d and power)
# ds : vector of Cohen’s d (e.g. c(0.2, 0.4, 0.6))
# target_powers: vector of target powers (e.g. c(0.80, 0.90))
run_all_eRTC <- function(n_sims = 5000,
mu_ctrl = 0,
ds = c(0.2, 0.4, 0.6),
target_powers = c(0.80, 0.90),
alpha = 0.05,
burn_in = 20,
ramp = 50,
c_max = 0.6,
seed = 42) {
set.seed(seed)
results <- list()
scenarios <- expand.grid(
hypothesized_d = ds,
target_power = target_powers,
stringsAsFactors = FALSE
all_results <- data.frame()
for (i in 1:nrow(scenarios)) {
hyp_d <- scenarios\$hypothesized_d[i]
tgt_pow <- scenarios\$target_power[i]
[…]
hyp_d, tgt_pow * 100))
[…]
# Type I error
[…]
t1 <- simulate_eRTC(
n_sims = n_sims,
mu_ctrl = mu_ctrl,
mu_trt = NULL, # Null true
hypothesized_d = hyp_d,
target_power = tgt_pow,
alpha = alpha,
burn_in = burn_in,
ramp = ramp,
c_max = c_max
# Power
[…]
pow <- simulate_eRTC(
n_sims = n_sims,
mu_ctrl = mu_ctrl,
mu_trt = mu_ctrl + hyp_d, # sd = 1, so delta = d
hypothesized_d = hyp_d,
target_power = tgt_pow,
alpha = alpha,
burn_in = burn_in,
ramp = ramp,
c_max = c_max
all_results <- rbind(all_results, data.frame(
hypothesized_d = hyp_d,
target_power = tgt_pow,
n_patients = t1\$n_patients,
type1_error = t1\$rejection_rate,
type1_se = t1\$se,
eRTC_power = pow\$rejection_rate,
power_se = pow\$se,
median_crossing = pow\$median_crossing
results\$summary <- all_results
[…]
print(all_results %>%
mutate(
hypothesized_d = sprintf("%.2f", hypothesized_d),
target_power = sprintf("%.0f%%", target_power * 100),
type1_error = sprintf("%.3f", type1_error),
eRTC_power = sprintf("%.1f%%", eRTC_power * 100)
select(hypothesized_d, target_power, n_patients,
type1_error, eRTC_power, median_crossing))
return(results)
# Example run (interactive)
if (interactive()) {
# Grid of d and power
results_eRTC <- run_all_eRTC(
n_sims = 2000, # lower for quick tests; bump to 5000+ for final
mu_ctrl = 0,
ds = c(0.2, 0.4, 0.6),
target_powers = c(0.80, 0.90),
alpha = 0.05,
burn_in = 20,
ramp = 50,
c_max = 0.6
# Trajectory plots - Alternative examples
p_alt_04_80 <- plot_trajectories_eRTC(
mu_ctrl = 0,
mu_trt = 0.4, # true d = 0.4
hypothesized_d = 0.4,
target_power = 0.80,
alpha = 0.05
ggsave("traj_eRTC_alt_d0.4_80pow.pdf", p_alt_04_80, width = 8, height = 5)
# Trajectory plots - Null
p_null_04_80 <- plot_trajectories_eRTC(
mu_ctrl = 0,
mu_trt = NULL, # null
hypothesized_d = 0.4,
target_power = 0.80,
alpha = 0.05
ggsave("traj_eRTC_null_d0.4_80pow.pdf", p_null_04_80, width = 8, height = 5)
# e-Survival: Sequential Log-Rank Test (Unified)
library(tidyverse)
library(survival) # Used for standard verification if needed
# Core e-Survival function
# time: vector of event times
# status: vector of status (1=event, 0=censored)
# treatment: vector of assignments (1=trt, 0=ctrl)
# burn_in: events to observe before betting
# ramp: number of events to ramp up betting intensity
# lambda: maximum betting scalar (0 < lambda < 1)
@@
burn_in = 20,
ramp = 50,
lambda_max = 0.2) {
@@
# 1. Create a dataframe and sort by time (events happen sequentially)
df <- data.frame(time, status, treatment) %>%
arrange(time)
# Vectors for processing
@@
# Track active Risk Set manually for speed/simplicity
# Initially, everyone is at risk (assuming time 0 start)
@@
# We need to process events one by one.
# Note: This simple loop assumes no tied times for clarity.
# (Ties are rare in continuous simulation).
@@
# --- A. Betting Step (Before observing outcome i) ---
# Determine bet direction based on history
@@
# Direction:
# If cumulative_Z is negative, Treatment has FEWER events than expected (Benefit).
# We bet that this trend continues.
# If Observed (1) - Expected (p) is likely negative, we want a NEGATIVE bet
# to make the wealth multiplier (1 + bet * U) > 1.
# Simple strategy: proportional to the sign of the current trend
@@
# Probability that *if* an event happens, it is in Treatment arm
# p_null = (Risk Treatment) / (Total Risk)
@@
if (total_risk > 0) {
p_null <- risk_trt / total_risk
} else {
p_null <- 0.5
# --- C. Update Wealth (Only on events, not censorings) ---
@@
# U_i: The "score" contribution (Observed - Expected)
# If Treatment event: 1 - p
# If Control event: 0 - p
@@
# Martingale update
# Constraint: ensure (1 + b_i * U_i) is non-negative.
# Since |U_i| < 1 and lambda_max < 1, this is generally safe.
@@
if (i > 1) {
wealth[i] <- wealth[i-1] * multiplier
@@
wealth[i] <- multiplier
} else {
# Censored event: No bet made, wealth carries forward
@@
# --- D. Update Risk Set for next step ---
# The person at index i is removed from risk set (event or censored)
@@
# Map wealth back to original unsorted indices is tricky in sequential plots,
# so we return the Sequential Wealth (wealth over event time).
return(data.frame(
event_num = 1:n,
time = T_sorted,
wealth = wealth
# Simulate a single survival trial (Weibull)
# HR < 1 implies Treatment is better (longer survival)
simulate_trial_survival <- function(n,
HR = 1,
shape = 1.2,
scale = 10,
cens_prop = 0.0) {
# 1. Treatments
@@
# 2. True Times (Weibull AFT)
# HR in Cox model corresponds roughly to scale change in Weibull
# Scale_trt = Scale_ctrl / (HR^(1/shape))
@@
# 3. Censoring
# Simple independent uniform censoring to achieve approx proportion
@@
C <- runif(n, 0, 2 * scale) # broad censoring distribution
# Adjust range to hit target prop roughly if needed, keeping simple here
@@
return(list(time = time, status = status, treatment = treatment))
# Unified simulation function
simulate_eSurvival <- function(n_sims = 1000,
n_patients = NULL, # Change to NULL to trigger auto-calc
HR_true = 1,
target_HR = 0.7,
target_power = 0.80, # Added for flexibility
alpha = 0.05,
burn_in = 30,
ramp = 50,
lambda_max = 0.25) {
# Determine simulation type
if (HR_true == 1) {
sim_type <- "TypeIError(Null)"
} else {
sim_type <- "Power(Alternative)"
# Calculate Sample Size if not provided
@@
# Total events needed
req_events <- ceiling( 4 * ((z_alpha + z_beta) / log(target_HR))^2 )
# Assuming no censoring, Patients = Events
n_patients <- req_events
[…]
target_power*100, target_HR, n_patients))
[…]

### v5

@@ 1 Introduction
Monitoring clinical trials, especially in the context of acutely ill patients, is of paramount importance. Traditional methods include $\alpha-spending$ functions. When interim analyses are spaced, evidence grows unnoticed between the interim analyses. This may both delay the implementation of potentially useful strategies or prolong trials when safety signals arise early.
@@ 4.2 Betting strategy for continuous outcomes
    A center and scale for the outcomes observed so far. Outcomes can be all over the place. We need to standardize.
@@ 4.6 Simulation overview
A visual representation is shown in Figure 3.
@@ 5.4 Simulation Results
6 Multi-State Models
Clinical trials in critical care increasingly use multi-state models to capture patient trajectories. A patient in the ICU may improve to a general ward, be discharged home, or die—and some of these transitions can reverse. Traditional analyses model the full transition matrix or use competing risks methods, requiring assumptions about the dependence structure between outcomes.
The e-RT framework extends naturally to this setting. Rather than modeling the Markov structure, we classify each transition as “good” (recovery-oriented) or “bad” (deterioration-oriented) and bet on this binary outcome. The method is agnostic to the underlying stochastic process: we simply ask whether good transitions predict treatment assignment.
6.1 Setup
Consider a four-state model common in ICU trials:
  • State 1: General Ward
  • State 2: ICU
  • State 3: Home (absorbing)
  • State 4: Dead (absorbing)
Patients begin in the ICU and are followed for a fixed period (e.g., 28 days). Each day, a patient may transition between states according to arm-specific transition probabilities.
We define good transitions as those representing clinical improvement:
  • ICU $\to$ Ward (step-down from intensive care)
  • Ward $\to$ Home (discharge)
All other transitions—including ICU $\to$ Dead, Ward $\to$ Dead, and Ward $\to$ ICU (readmission)—are classified as bad.
The null hypothesis is that the rate of good transitions is equal between arms:
$H_{0}:P(\text{good transition}\mid\text{treatment})=P(\text{good transition}\mid\text{control})$ (23)
6.2 Betting Strategy
Each state transition generates one betting opportunity. Unlike the binary e-RT where each patient contributes one observation, here a single patient may contribute multiple transitions as they move through states. A patient who goes ICU $\to$ Ward $\to$ ICU $\to$ Ward $\to$ Home contributes four transitions and four bets.
Let $\hat{\delta}_{i-1}$ be the difference in good-transition rates between arms, estimated from all transitions observed before transition $i$:
[…]
The betting fraction follows the same logic as binary e-RT:
[…]
where $c_{i}$ ramps from 0 to 1 over a burn-in period, exactly as before.
The wealth update remains:
$W_{i}=W_{i-1}\times\begin{cases}\lambda_{i}/0.5&\text{if treatment arm}\\ (1-\lambda_{i})/0.5&\text{if control arm}\end{cases}$ (26)
6.3 Validity
The martingale property holds by the same argument as before. Under the null, treatment assignment is independent of transition quality. Given any transition (good or bad), the probability it came from the treatment arm is 0.5. The expected wealth multiplier is therefore 1, and $(W_{i})$ is a test martingale.
The key insight is that validity does not depend on the Markov assumption or any specific transition structure. We have simply replaced “patient with event” with “transition classified as good” and applied the same betting logic. The data-generating process happens to be Markovian in our simulations, but the test would remain valid for any process that generates a sequence of binary-classified transitions from randomized patients.
6.4 Simulation Study
We simulated trials with the following daily transition probabilities:
Table 4: Daily Transition Probabilities
To State
From State Ward ICU Home Dead
Control
Ward 0.880 0.070 0.030 0.020
ICU 0.070 0.915 0.000 0.015
Treatment
Ward 0.870 0.050 0.050 0.030
ICU 0.090 0.900 0.000 0.010
Treatment improves recovery transitions: ICU $\to$ Ward increases from 7% to 9% daily, and Ward $\to$ Home increases from 3% to 5% daily. The effect on mortality is modest.
All patients started in the ICU and were followed for 28 days. At day 28, the expected state distributions were:
Table 5: Day 28 State Distribution
Arm Ward ICU Home Dead
Control 20.8% 26.3% 18.8% 34.1%
Treatment 16.9% 16.6% 33.5% 33.0%
Difference $3.9\%$ $9.7\%$ $4.7\% 1.1$$1\%$
Treatment substantially increases discharge home ($14.7\%$) while barely changing mortality ($1.1\%$). This is precisely the scenario where trajectory-based analysis adds value: a mortality-only endpoint would detect little effect, while multi-state analysis captures the clinically meaningful improvement in recovery.
We used burn-in = 30 transitions, ramp = 50 transitions, and threshold $1/\alpha=20$. Results from 1,000 simulated trials at $N=1000$ patients:
Table 6: Operating Characteristics for e-RTms ($N=1000$)
Metric Value
Type I Error 0.30% (SE: 0.17%)
Power 89.3% (SE: 1.0%)
Median transitions per trial 2017
Median crossing 898 transitions (45%)
Type I error was well below the nominal 5%, consistent with the conservative behavior observed in other e-RT variants. Power was 89% at $N=1000$; sample size search indicated approximately $N=800$ for 80% power.
The median trial generated approximately 2000 transitions from 1000 patients (roughly 2 transitions per patient). When the threshold was crossed, it occurred at 898 transitions—45% of the way through the trial.
6.5 Trajectory Examples
Figure 5 shows 30 example wealth trajectories under the null and alternative hypotheses. Under the null (left panel), wealth fluctuates around 1 with the characteristic downward drift from betting on noise. Under the alternative (right panel), wealth grows systematically as good transitions accumulate preferentially in the treatment arm.
Figure 5: Trajectories of e-RTms for a trial with $N=1000$ patients. Left: under the null hypothesis (equal transition matrices), wealth fluctuates randomly. Right: under the alternative hypothesis, wealth grows as treatment improves recovery transitions. Dashed red line: rejection threshold ($1/\alpha=20$).
6.6 Interpretation
The e-RTms approach offers a philosophically clean handling of multi-state outcomes. By collapsing the transition matrix into a single binary classification, we sidestep the complexities of competing risks and counterfactual reasoning. We do not ask “what would have happened to this patient had they been in the other arm?” We simply ask “does knowing whether this transition was good help predict which arm the patient was in?”
This comes at a cost: we lose information about which specific transitions drive the effect. A treatment that dramatically improves ICU $\to$ Ward transitions but worsens Ward $\to$ Home transitions might show no overall effect if the good-transition rates balance. For monitoring purposes, this limitation is acceptable—the goal is to detect any departure from exchangeability, not to dissect the mechanism. Detailed transition-specific analyses can follow at trial completion. e-RT is a monitoring, not a reporting, method.
The method is most valuable when treatment affects trajectory without substantially changing mortality— the pattern in our simulation. For trials where mortality is the dominant signal, e-survival may be more powerful. For trials where the clinical question is “does treatment help patients recover faster?”, e-RTms provides a natural framework.
@@ 7 Discussion
The e-RT is a nonparametric sequential test for randomized trials based on the betting framework for e-values (i-bet pmlr-v177-duan22a). The method requires only that treatment assignment is randomized — no distributional assumptions about outcomes are needed. This makes it a robust complement to model-based analyses. It can be extended to continuous endpoints and time-to-event analyses.
@@ 8.4 Version Control
    Fourth Version (Dec 17, 2025): Correction on signal direction on e-RTC. Text adjustments
  5. 5.
    Fifth Version (Dec 31, 2025): Added multi-state extension (e-RTms).
@@ Appendix A R Code
# e-RT: Sequential Randomization Tests Using e-values
# Supplementary R Code
@@
lambda <- if (outcome[i] == 1) {
0.5 + 0.5 * c_i * delta_hat
@@
multiplier <- if (treatment[i] == 1) lambda / p else (1 - lambda) / (1 - p)
@@
data.frame(treatment = treatment, outcome = outcome)
simulate_eRT <- function(n_sims = 5000, p_ctrl, p_trt = NULL,
hypothesized_ARR = NULL, target_power = 0.80,
alpha = 0.05, burn_in = 50, ramp = 100) {
@@
if (is.null(hypothesized_ARR)) hypothesized_ARR <- p_ctrl - p_trt
@@
wealth <- compute_eRT(trial\$treatment, trial\$outcome,
burn_in = burn_in, ramp = ramp)
@@
list(
@@
median_crossing = median(first_crossing, na.rm = TRUE),
@@
plot_trajectories <- function(n_trials = 30, p_ctrl, p_trt = NULL,
hypothesized_ARR = NULL, target_power = 0.80,
alpha = 0.05, burn_in = 50, ramp = 100,
@@
if (is.null(hypothesized_ARR)) hypothesized_ARR <- p_ctrl - p_trt
if (is.null(p_trt)) p_trt <- p_ctrl
@@
title <- if (true_ARR == 0) {
sprintf("e-RTUnderNull(n=%d,designedfor%.0f%%ARR)",
@@
sprintf("e-RTUnderAlternative(n=%d,%.0f%%power)",
@@
trajectories <- lapply(1:n_trials, function(i) {
@@
wealth <- compute_eRT(trial\$treatment, trial\$outcome,
burn_in = burn_in, ramp = ramp)
data.frame(patient = 1:n_patients, wealth = wealth, trial = i)
@@
ggplot(df, aes(x = patient, y = wealth, group = trial)) +
@@
labs(title = title,
@@
x = "Patient", y = "e-value(logscale)") +
@@
theme(plot.title = element_text(face = "bold"), panel.grid.minor = element_blank())
# --- e-RTC: Continuous Outcomes ---
compute_eRTC <- function(treatment, outcome, p = 0.5, burn_in = 20,
ramp = 50, c_max = 0.6) {
@@
g_i <- s_i / (1 + abs(s_i))
@@
direction <- if (is.nan(mean_trt) $|$ is.nan(mean_ctrl)) 0 else sign(mean_trt - mean_ctrl)
@@
multiplier <- if (treatment[i] == 1) lambda / p else (1 - lambda) / (1 - p)
@@
simulate_trial_continuous <- function(n, p_trt = 0.5, mu_ctrl = 0, mu_trt = 0, sd = 1) {
@@
data.frame(treatment = treatment, outcome = outcome)
simulate_eRTC <- function(n_sims = 5000, mu_ctrl = 0, mu_trt = NULL,
hypothesized_d, target_power = 0.80, alpha = 0.05,
burn_in = 20, ramp = 50, c_max = 0.6) {
if (missing(hypothesized_d)) stop("Mustspecifyhypothesized_d")
@@
$s$ <- power.t.test(delta = delta_design, sd = sd_true, power = target_power,
sig.level = alpha, type = "two.sample", alternative = "two.sided")
@@
trial <- simulate_trial_continuous(n_patients, mu_ctrl = mu_ctrl,
mu_trt = mu_trt, sd = sd_true)
wealth <- compute_eRTC(trial\$treatment, trial\$outcome, p = 0.5,
burn_in = burn_in, ramp = ramp, c_max = c_max)
@@
list(
@@
median_crossing = median(first_crossing, na.rm = TRUE),
@@
plot_trajectories_eRTC <- function(n_trials = 30, mu_ctrl = 0, mu_trt = NULL,
hypothesized_d, target_power = 0.80,
alpha = 0.05, burn_in = 20, ramp = 50,
c_max = 0.6, title = NULL) {
if (missing(hypothesized_d)) stop("Mustspecifyhypothesized_d")
@@
$s$ <- power.t.test(delta = hypothesized_d, sd = sd_true, power = target_power,
sig.level = alpha, type = "two.sample", alternative = "two.sided")
@@
title <- if (true_d == 0) {
sprintf("e-RTCUnderNull(n=%d,designedford=%.2f)", n_patients, hypothesized_d)
@@
sprintf("e-RTCUnderAlternative(n=%d,trued=%.2f)", n_patients, true_d)
trajectories <- lapply(1:n_trials, function(i) {
trial <- simulate_trial_continuous(n_patients, mu_ctrl = mu_ctrl,
mu_trt = mu_trt, sd = sd_true)
wealth <- compute_eRTC(trial\$treatment, trial\$outcome, p = 0.5,
burn_in = burn_in, ramp = ramp, c_max = c_max)
data.frame(patient = 1:n_patients, wealth = wealth, trial = i)
@@
ggplot(df, aes(x = patient, y = wealth, group = trial)) +
@@
labs(title = title,
@@
x = "Patient", y = "e-value(logscale)") +
annotate("text", x = n_patients * 0.95, y = (1/alpha) * 1.5,
label = sprintf("1/alpha=%.0f", 1/alpha), color = "red", hjust = 1) +
@@
theme(plot.title = element_text(face = "bold"), panel.grid.minor = element_blank())
# --- e-Survival: Time-to-Event Outcomes ---
library(survival)
@@
burn_in = 20, ramp = 50, lambda_max = 0.2) {
@@
df <- data.frame(time, status, treatment) %>% arrange(time)
@@
p_null <- if (total_risk > 0) risk_trt / total_risk else 0.5
@@
wealth[i] <- if (i > 1) wealth[i-1] * multiplier else multiplier
@@
data.frame(event_num = 1:n, time = T_sorted, wealth = wealth)
simulate_trial_survival <- function(n, HR = 1, shape = 1.2, scale = 10, cens_prop = 0.0) {
@@
C <- runif(n, 0, 2 * scale)
@@
list(time = time, status = status, treatment = treatment)
simulate_eSurvival <- function(n_sims = 1000, n_patients = NULL, HR_true = 1,
target_HR = 0.7, target_power = 0.80, alpha = 0.05,
burn_in = 30, ramp = 50, lambda_max = 0.25) {
@@
n_patients <- ceiling(4 * ((z_alpha + z_beta) / log(target_HR))^2)